The speed of sound where a tuning fork of frequency 262 Hz produces the third resonance position above a closed air column that is 1.59 m in length is ___m/s. (Give your answer with the correct number of sig digs and do not include units in your answer)


1
Expert's answer
2021-07-22T09:55:52-0400

The third resonance position corresponds to the fifth harmonic of the closed air column. Therefore, we can find the wavelength of the wave:


"\lambda=\dfrac{4}{5}L=\dfrac{4}{5}\cdot1.59\ m=1.272\ m."

Finally, we can find the speed of sound:


"v=f\lambda,""v=262\ Hz\cdot1.272\ Hz=333\ \dfrac{m}{s}."